OpenStudy (anonymous):

Find 4x − 2y if x = −1 and y = −3

OpenStudy (anonymous):

Just plug in -1 for x and -3 for y and solve

OpenStudy (anonymous):

Substitute the value of x and y into the expression

OpenStudy (anonymous):

4(-1) - 2(-3) = -4 - (-6) = -4 + 6 = 2. Full solution for you!

OpenStudy (anonymous):

so you get 4(-1)-2(-3), then solve

OpenStudy (anonymous):

4x − 2y = 4(-1) - 2(-3) = −4 + 6 = 2

OpenStudy (anonymous):

-4+6 =2

OpenStudy (anonymous):

@kkj correct

OpenStudy (anonymous):

Thank you for the appreciation